# Robust Optimization with Diffusion Models for Green Security

### v2

@@ 3 Preliminaries on Diffusion Model
[…]
@@ 5.1 Mixed strategy over mixed strategies
[…]
@@
In this reformulation, the adversary’s pure strategy is no longer a single value but instead a full distribution $\tau(\mathbf{z})$. Consequently, the adversary’s pure strategy space becomes $\mathcal{T}$ and the corresponding mixed strategy space is the set of distributions over these distributions, $\Delta(\mathcal{T})$. Under this framework, the defender’s utility function takes the form $\mathbb{E}_{\tau(\mathbf{z})}\bigl{[}u(\mathbf{x},\mathbf{z})\bigr{]}$, while the attacker’s utility becomes $-\mathbb{E}_{\tau(\mathbf{z})}\bigl{[}u(\mathbf{x},\mathbf{z})\bigr{]}$.
@@ 5.2 Double Oracle Flow
$\mathbf{x}_{i}=\arg\max_{\mathbf{x}\in\mathcal{X}}U(\mathbf{x},\hat{\sigma}^{*}_{i-1}).$ (5)
@@
3:$\mathbf{x}_{0}\leftarrow\text{random strategy}$,  $\tau_{0}\leftarrow p_{\theta}(\mathbf{z}\mid\mathbf{c})$
4:$\mathcal{X}_{0}\leftarrow\{\mathbf{x}_{0}\}$,  $\mathcal{T}_{0}\leftarrow\{\tau_{0}\}$
@@
7:    $\hat{\mathcal{T}}_{i-1}\leftarrow\text{Empirical distributions from }\mathcal{T}_{i-1}$ using Alg. 2
@@ 5.3 Sampling with Twisted Sequential Monte Carlo
5:$\quad\quad\{\mathbf{z}_{n}^{t}\}_{n=1}^{N}\sim\mathrm{Multinomial}\bigl{(}\{\mathbf{z}_{n}^{t}\}_{n=1}^{N};\,\{w_{n}^{t}\}_{n=1}^{N}\bigr{)}$
@@
$U(\mathbf{x},\hat{\tau}(\mathbf{z}))\to U(\mathbf{x},\tau(\mathbf{z}))\quad\text{almost surely},$
@@ 5.4 Convergence Analysis
In practice, the attacker’s action at each target must lie in a bounded interval, e.g. $[0,z_{\mathrm{max}}]$. For instance, the number of snares at any region cannot exceed a practical upper limit. Consequently, it is reasonable to treat the action space as compact, ensuring that $p_{\theta}(\mathbf{z}\mid\mathbf{c})$ has compact support.
@@
  • Item 1: Without terminating condition, every weakly convergent subsequence of Alg. 1 converges to an $\epsilon$-equilibrium in a possibly infinite number of iterations. Such a weakly convergent subsequence always exist.
  • Item 2: With the terminating condition, Alg. 1 terminates in a finite number of iterations. Also, it converges to a finitely supported $4\epsilon$-equilibrium with probability at least $1-\mathsf{prob}$.
@@
  • Step 2: We show that, under our finite sampling scheme, the probability that the maximum cell-wise error exceeds $\epsilon/4$ is nonzero only during the first $i_{r}$ iterations, for some finite $i_{r}$.
  • Step 3: We treat the strategies generated in the first $i_{r}$ rounds as the initial strategy set in the standard Double Oracle (DO) algorithm [Adam et al., 2021]. We then adapt the original convergence proof to account for the error introduced by finite sampling, which is now bounded by $\epsilon/4$.
By relaxing the error bound in Item 1, we obtain convergence within a finite number of iterations. The additional approximation error in Item 2 stems from two sources: (1) enforcing finite termination, and (2) using estimated utilities of mixed strategy pairs when evaluating the stopping condition.
@@ Appendix A More Details on Diffusion Models
[…]
@@ Appendix B Examples of mixed strategy over mixed strategies
$U(\mathbf{x},\hat{\tau}(\mathbf{z}))\to U(\mathbf{x},\tau(\mathbf{z}))\quad\text{almost surely},$
@@ Appendix F Definition of weak convergence
$|U(\pi_{i},\sigma_{i})-U(\pi_{i},\hat{\sigma}_{i})|=\sum_{\mathbf{x}\in\mathcal{X}_{i}}\sum_{\tau\in\mathcal{T}_{i}}\pi_{i}(\mathbf{x})\cdot\sigma_{i}(\tau)\cdot|U(\mathbf{x},\tau)-U(\mathbf{x},\hat{\tau})|$ (8)
$|U(\mathbf{x},\tau)-U(\mathbf{x},\hat{\tau})|$ denotes the sample estimation error for the pure strategy pair $(\mathbf{x},\tau)$ in the payoff matrix. The maximum on the right-hand side of Equation 8 is obtained when putting all the probability mass on the strategy pair with the largest sample estimation error, which is $\Delta_{i}$.
@@
Because $\delta>0$, $\sum_{i=1}^{\infty}\frac{1}{i^{1+\delta}}$ is a convergent series. Therefore, $\sum_{i=1}^{\infty}P(A_{i})<\infty$. From Borel-Cantelli Lemma, we then obtain $P(\limsup\limits_{i\to\infty}A_{i})=0$, which implies $A_{i}$ only happens for finite times. There exists $i_{r}$ that for any $i>i_{r}$,
$P(\Delta_{i}\geq\frac{\epsilon}{4})=0.$
@@
From the proof above, $A_{i}$ only happens for finite times. Assume $i_{r}$ is the largest number satisfying that $A_{i}$ happens. We then treat $(\mathcal{X}_{i_{r}},\mathcal{T}_{i_{r}})$ as the initial set of strategies for both players. Then our sampling scheme ensures that for any strategy pair $(\pi,\sigma)$ and iteration $i$, we have $|U(\pi_{i},\sigma_{i})-U(\pi_{i},\hat{\sigma}_{i})|\leq\Delta_{i}\leq\epsilon/4$.
@@
We then show that adding the terminating condition, for any $\epsilon>0$, algorithm 1 can terminate in a finite number of iterations. Also, when it stops, it converges to a $4\epsilon$-equilibrium with high probability.
@@
Consider now an infinite game, from Item 1 in Theorem 5.2, we know that $\bar{v}_{i}-\underline{v}_{i}\rightarrow\epsilon$. Also, our sampling scheme ensures that for any strategy pair $(\pi,\sigma)$ and iteration $i$ after some finite rounds $i_{r}$, we have $|U(\pi_{i},\sigma_{i})-U(\pi_{i},\hat{\sigma_{i}})|\leq\Delta_{i}\leq\frac{\epsilon}{4}$. This indicates that with $\epsilon>0$, the terminating condition will be satisfied after a finite number of iterations. Assume that the algorithm ends at the $j$-th iteration. This implies

### v3

@@ 3 Preliminaries on Diffusion Model
Here, $q(\mathbf{z}^{t})=\int q(\mathbf{z}^{0})\,q(\mathbf{z}^{t}\mid\mathbf{z}^{0})\,d\mathbf{z}^{0}$, and the gradient $\nabla_{\mathbf{z}^{t}}\!\log q(\mathbf{z}^{t})$ points toward regions of higher data density. In practice, we do not know $q(\mathbf{z}^{t})$ in closed form, so a neural score network $s_{\theta}(\mathbf{z}^{t},t)$ is trained to approximate this gradient via denoising score matching [Vincent, 2011, Ho et al., 2020]. Consequently, the learned reverse (denoising) transition becomes
@@ 5.1 Mixed strategy over mixed strategies
Let $\mathcal{T}$ denote the space of mixed strategies, where each $\tau\in\mathcal{T}$ represents a probability distribution over pure strategies. A mixed strategy over mixed strategies, $\sigma$, is a probability distribution over $\mathcal{T}$, formally expressed as $\sigma\in\Delta(\mathcal{T})$. This implies that $\sigma$ satisfies the following conditions: (1) $\sigma(\tau)\geq 0$ for all $\tau\in\Delta$, and (2) $\int_{\mathcal{T}}\sigma(\tau)\,d\tau=1$.
@@
In this reformulation, the adversary’s pure strategy is no longer a single value but instead a full distribution $\tau(\mathbf{z})$. Consequently, the adversary’s pure strategy space becomes $\mathcal{T}$ and the corresponding mixed strategy space is the set of distributions over these distributions, $\Delta(\mathcal{T})$. Under this framework, the defender’s utility function takes the form $\mathbb{E}_{\tau(\mathbf{z})}\bigl[u(\mathbf{x},\mathbf{z})\bigr]$, while the attacker’s utility becomes $\mathbb{E}_{\tau(\mathbf{z})}\bigl[u(\mathbf{x},\mathbf{z})\bigr]$.
@@ 5.2 Double Oracle Flow
$\vskip-10.00002pt\mathbf{x}_{i}=\arg\max_{\mathbf{x}\in\mathcal{X}}U(\mathbf{x},\hat{\sigma}^{*}_{i-1}).\vskip-10.00002pt$ (5)
@@
[…]
@@
[…]
@@ 5.3 Sampling with Twisted Sequential Monte Carlo
[…]
@@
[…]
@@ 5.4 Convergence Analysis
In practice, the attacker’s action at each target must lie in a bounded interval, e.g. $0,z_{\mathrm{max}}]$. For instance, the number of snares at any region cannot exceed a practical upper limit. Consequently, it is reasonable to treat the action space as compact, ensuring that $p_{\theta}(\mathbf{z}\mid\mathbf{c})$ has compact support.
@@
  • Item 1: Without terminating condition, every weakly convergent subsequence of Alg. 1 converges to an $\epsilon$-equilibrium in a possibly infinite number of iterations almost surely. Such a weakly convergent subsequence always exist.
  • Item 2: With the terminating condition, Alg. 1 terminates in a finite number of iterations almost surely. Also, it converges to a finitely supported $4\epsilon$-equilibrium with probability at least $1-\mathsf{prob}$.
@@
  • Step 2: We show that, under our finite sampling scheme, almost surely, the realized sequence of payoff matrices has a finite cutoff index after which all subsequent iterations have maximum cell-wise errors below $\epsilon/4$.
  • Step 3: We treat the strategies generated during the finite initial phase (before the cutoff index ensured by Step 2) as the initial strategy set in the standard Double Oracle (DO) algorithm [Adam et al., 2021]. We then adapt the original convergence proof to account for the error introduced by finite sampling, which is now bounded by $\epsilon/4$.
By relaxing the error bound in Item 1, we obtain convergence within a finite number of iterations almost surely. The additional approximation error in Item 2 stems from two sources: (1) enforcing finite termination, and (2) using estimated utilities of mixed strategy pairs when evaluating the stopping condition.
@@ Appendix A More Details on Diffusion Models
Here, $q(\mathbf{z}^{t})=\int q(\mathbf{z}^{0})\,q(\mathbf{z}^{t}\mid\mathbf{z}^{0})\,d\mathbf{z}^{0}$, and the gradient $\nabla_{\mathbf{z}^{t}}\!\log q(\mathbf{z}^{t})$ points toward regions of higher data density. In practice, we do not know $q(\mathbf{z}^{t})$ in closed form, so a neural score network $s_{\theta}(\mathbf{z}^{t},t)$ is trained to approximate this gradient via denoising score matching [Vincent, 2011, Ho et al., 2020]. Consequently, the learned reverse (denoising) transition becomes
@@ Appendix B Examples of mixed strategy over mixed strategies
[…]
@@ Appendix F Definition of weak convergence
[…]
@@
Because $\delta>0$, $\sum_{i=1}^{\infty}\frac{1}{i^{1+\delta}}$ is a convergent series. Therefore, $\sum_{i=1}^{\infty}P(A_{i})<\infty$. From Borel-Cantelli Lemma, we then obtain $P(\limsup\limits_{i\to\infty}A_{i})=0$. Let’s use $\Omega$ to denote the space of all payoff matrix sequence across iterations of Alg. 1 without terminating condition, and use $\omega\in\Omega$ to denote a realization. $P(\limsup\limits_{i\to\infty}A_{i})=0$ implies that
[…]
Intuitively, this means almost surely, the realized sequence of payoff matrices has a finite cutoff index after which all estimation errors remain below $\epsilon/4$.
@@
Recall with probability $1$, the realized sequence of payoff matrices has a finite cutoff index after which all estimation errors remain below $\epsilon/4$. From now on, we fix any such realization $\omega$ from this probability $1$ set, and omit the notational dependence on $\omega$ for simplicity. From the proof above, $A_{i}$ only happens for finite times. Assume $i_{r}$ is the largest number satisfying that $A_{i}$ happens. We then treat $(\mathcal{X}_{i_{r}},\mathcal{T}_{i_{r}})$ as the initial set of strategies for both players. Then our sampling scheme ensures that for any strategy pair $(\pi,\sigma)$ and iteration $i$, we have $|U(\pi_{i},\sigma_{i})-U(\pi_{i},\hat{\sigma}_{i})|\leq\Delta_{i}\leq\epsilon/4$.
@@
We then show that adding the terminating condition, for any $\epsilon>0$, algorithm 1 can terminate in a finite number of iterations almost surely. Also, when it stops, it converges to a $4\epsilon$-equilibrium with high probability.
@@
Consider now an infinite game, from Item 1 in Theorem 5.2, we know that $\bar{v}_{i}-\underline{v}_{i}\rightarrow\epsilon$. Also, our sampling scheme ensures that for any strategy pair $(\pi,\sigma)$, almost surely, there exists a finite cutoff index after which all subsequent iteration $i$ satisfy $|U(\pi_{i},\sigma_{i})-U(\pi_{i},\hat{\sigma_{i}})|\leq\Delta_{i}\leq\frac{\epsilon}{4}$. This indicates that with $\epsilon>0$, the terminating condition will be satisfied after a finite number of iterations almost surely. Assume that the algorithm ends at the $j$-th iteration. This implies